A man buys a house for $200,000. He makes a $50,000 down payment and agrees to

amortize the rest of the debt with quarterly payments over the next 10 years. If the

interest on the debt is 12%, compounded quarterly, find

(i) Size of the quarterly payments, (4 marks)

(ii) Total amount of the payments, and (4 marks)

(iii) Total amount of interest paid.


1
Expert's answer
2021-10-04T16:19:30-0400

i)

The house is 200000. The down payment is 50000, bringing the value of the loan to 150000.

So:

"P=L\frac{c(1+c)^n}{(1+c)^n-1}=150000\cdot\frac{0.3(1+0.3)^{40}}{(1+0.3)^{40}-1}=\$6489.36"


ii)

"6489.36\cdot40=\$259574.27"


iii)

"259574.27-150000=\$109574.27"